# Distributed High-Dimensional Quantile Regression: Estimation Efficiency and Support Recovery

### v1

@@ 2.3 Distributed Estimation with a Double-smoothing Shifted Loss Function
Suppose the random sample ${\cal Z}^{N}=\{(\boldsymbol{X}_{i},Y_{i})\}_{i=1}^{N}$ are randomly stored in $m$ machines $\mathcal{M}_{1},\ldots,\mathcal{M}_{m}$ with the equal local sample size that $n=N/m$. Without loss of generality, we assume that $\mathcal{M}_{1}$ is the central machine and denote those samples in the $k$-th machine as $\{(\boldsymbol{X}_{i},Y_{i})\}_{i\in\mathcal{M}_{k}}$ with $|\mathcal{M}_{k}|=n$, for $k=1,\ldots,m$. Based on the initial estimator $\widehat{\boldsymbol{\beta}}_{0}$, every local machines can compute the transformed samples as $\{(\widetilde{\boldsymbol{X}}_{i,h},\widetilde{Y}_{i,h})\}_{i\in\mathcal{M}_{k}}$ according to (6) and (7). For ease of notation, let
@@
$\widehat{\boldsymbol{D}}_{h}=\frac{1}{m}\sum_{k=1}^{m}\widehat{\boldsymbol{D}}_{k,h}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}}$
@@
From the Taylor expansion of $\mathcal{L}_{N}(\boldsymbol{\beta})$ around $\widehat{\boldsymbol{\beta}}_{0}$, we have
@@
It is easy to see that $\partial\mathcal{L}_{N}(\widehat{\boldsymbol{\beta}}_{0})$ and $\widehat{\boldsymbol{D}}_{h}$ can be simply calculated by averaging the local ones. However, the burden of transmitting the local $p\times p$ covariance matrix $\widehat{\boldsymbol{D}}_{k,h}$ is heavy when $p$ is large. To save the communication cost, we replace the global Hessian $\widehat{\boldsymbol{D}}_{h}$ with the local Hessian $\widehat{\boldsymbol{D}}_{1,b}$, here $h$ and $b$ denote the global bandwidth and local bandwidth, respectively, and we assume $b\geq h\geq 0$. So we can rewrite (11) as
$\mathcal{L}_{N}(\boldsymbol{\beta};\widehat{\boldsymbol{D}}_{h})$ $=\underbrace{\mathcal{L}_{N}(\boldsymbol{\beta};\widehat{\boldsymbol{D}}_{1,b})}_{(i)\text{ Shifted loss}}$ (12)
$+\underbrace{\mathcal{O}_{P}\left\{\|\widehat{\boldsymbol{D}}_{h}-\widehat{\boldsymbol{D}}_{1,b}\|_{op}\cdot|\boldsymbol{\beta}-\widehat{\boldsymbol{\beta}}_{0}|_{2}^{2}\right\}}_{(ii)\text{Approximation error}}.$
Note that the substituted local Hessian matrix is flexibly controlled by a local bandwidth $b$ instead of the global bandwidth $h$, which ensures that $\|\widehat{\boldsymbol{D}}_{h}-\widehat{\boldsymbol{D}}_{1,b}\|_{op}=\mathcal{O}_{\mathbb{P}}(1)$ (Detailed proof can be referred in Appendix A.3). Remove the terms that are independent of $\boldsymbol{\beta}$ in (i) and the negligible approximation error (ii) in (12), the shifted loss function can be simplified to
@@
$-\boldsymbol{\beta}^{\mathrm{T}}\left\{\boldsymbol{z}_{N}+(\widehat{\boldsymbol{D}}_{1,b}-\widehat{\boldsymbol{D}}_{h})\widehat{\boldsymbol{\beta}}_{0}\right\},$
where $z_{N}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{Y}_{i,h}$. Up to now, we only need to focus on the shifted loss function in (13) instead of the pseudo global loss function in (10). Specifically, we define the one-step distributed estimator as
@@
In particular, local machines only need to compute two vectors $z_{n,k}=\frac{1}{n}\sum_{i\in\mathcal{M}_{k}}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{Y}_{i,h}$ and $\widehat{\boldsymbol{D}}_{k,h}\widehat{\boldsymbol{\beta}}_{0}=\frac{1}{n}\sum_{i\in\mathcal{M}_{k}}\widetilde{\boldsymbol{X}}_{i,h}(\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{0})$ and then broadcast them to the central machine with communication cost of $\mathcal{O}(mp)$. The central machine firstly calculates $z_{N}$ and $\widehat{\boldsymbol{D}}_{h}\widehat{\boldsymbol{\beta}}_{0}$ by simple averaging, then solves (14) via some well-learned algorithms, e.g., the PSSsp algorithm (Schmidt, 2010), the active set algorithm (Solntsev et al., 2015) and the coordinate descent algorithm (Wright, 2015).
Given $\widehat{\boldsymbol{\beta}}_{1,h}$ as the estimator from the first iteration, we construct an iterative distributed estimation procedure. Specifically, in the $t$-th iteration, we update the pseudo covariates and responses in (6) and (7) by substituting $\widehat{\boldsymbol{\beta}}_{0}$ with $\widehat{\boldsymbol{\beta}}_{t-1,h}$,
@@
for $i=1,\ldots,N$, where $\widehat{e}^{(t)}_{i,h}=Y_{i}-\boldsymbol{X}_{i}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{t,h}$. Similar with (14), our iterative DHSQR estimator is
@@
Algorithm 1 Distributed high-dimensional sparse quantile regression (DHSQR)
@@
$(\widehat{\boldsymbol{D}}_{1,b}^{(t)}-\widehat{\boldsymbol{D}}_{h}^{(t)})\widehat{\boldsymbol{\beta}}_{t-1,h}=\widehat{\boldsymbol{D}}_{1,b}^{(t)}\widehat{\boldsymbol{\beta}}_{t-1,h}-\frac{1}{m}\sum_{k=1}^{m}\widehat{\boldsymbol{D}}_{k,h}^{(t)}\widehat{\boldsymbol{\beta}}_{t-1,h},\quad\boldsymbol{z}_{N}^{(t)}=\frac{1}{m}\sum_{k=1}^{m}\boldsymbol{z}_{n,k}^{(t)}$
@@
In this paper, we adopt the coordinate descent algorithm to solve (17). For the choice of the initial estimator $\widehat{\beta}_{0}$, we take the solution of $\ell_{1}$-penalized QR regression using the local data on the central machine, which can be solved by the R package ”quantreg(Koenker, 2005). Other types of initialization are available as long as it satisfies Assumption 3.6 in Section 3. We summarize the entire distributed estimation procedure in Algorithm 1.
Space complexity. In each local machine, DHSQR method necessitates storing $\boldsymbol{X}_{i}\in\mathbb{R}^{(p+1)\times n}$, $Y_{i}\in\mathbb{R}^{n}$, $\tilde{\boldsymbol{X}}_{i,h}^{(t)}\in\mathbb{R}^{(p+1)\times n}$, and $\tilde{Y}_{i,h}^{(t)}\in\mathbb{R}^{n}$, resulting in a space complexity of order $\mathcal{O}(np)$. Additionally, to solve the lasso problem, we require storing a $(p+1)\times(p+1)$ matrix, resulting in a space complexity not exceeding $\mathcal{O}\left(p^{2}\right)$. Hence, the overall space complexity is of order $\mathcal{O}\left(np+p^{2}\right)$ for each local machine. The total space complexity of the total system sums up to $\mathcal{O}\left(Np+mp^{2}\right)$.
@@ 3 Statistical Guarantees
$S=\{j:\beta_{j}\neq 0,j\in\mathbb{N}_{+}\}\subseteq\{1,\ldots,p\}$
@@
Assume that the kernel function $K(\cdot)$ is symmetric, non-negative, bound and integrates to one. In addition, the kernel function satisfies that $\int_{-\infty}^{\infty}u^{2}K(u)\mathrm{d}u<\infty$ and $\min_{|u|\leq 1}K(u)>0$.
@@
There exist $f_{2}\geq f_{1}>0$ such that $f_{1}\leqslant f_{\varepsilon\mid\boldsymbol{X}}(0)\leqslant f_{2}$ almost surely (for all $\boldsymbol{X}$). Moreover, there exists some $l_{0}$ such that $\left|f_{\varepsilon\mid\boldsymbol{X}}(u)-f_{\varepsilon\mid\boldsymbol{X}}(v)\right|\leq l_{0}|u-v|$ for any $u,v\in\mathbb{R}$ and all $\boldsymbol{X}$.
@@
The random covariate $\boldsymbol{X}\in\mathbb{R}^{p}$ is sub-Gaussian: there exists some $c_{1}>0$ such that ${\mathbb{P}}\left(\left|\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\Sigma}^{-1/2}\delta\right|\geqslant c_{1}t\right)\leq 2e^{-t^{2}/2}$ for every unit vector $\delta$ and $t>0$, where $\boldsymbol{\Sigma}=\mathbb{E}(\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}})$.
@@
Denote $\boldsymbol{I}=\mathbb{E}\left\{f_{\varepsilon\mid\boldsymbol{X}}(0)\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}\right\}$, then $\boldsymbol{I}$ satisfies that $\left\|\boldsymbol{I}_{S^{c}\times S}\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\leq 1-\alpha,$ for some $10<\alpha<$ 1. Moreover, we assume that $\lambda_{-}\leq\Lambda_{\min}(\boldsymbol{I})\leq\Lambda_{\max}(\boldsymbol{I})\leq\lambda_{+}$ for some $\lambda_{-},\lambda_{+}>0$.
@@
The dimension $p$ satisfies $p=O\left(N^{\nu}\right)$ for some $\nu>0$. The local sample size $n$ satisfies $n\geq N^{c}$ for some $10<c<$ 1, and the sparsity level $s$ satisfies $s=O\left(n^{r}\right)$ for some $30\leq r<1/$ 3.
@@
Assumption 3.1 imposes some regularity conditions on the kernel function $K(\cdot)$, which is satisfied by many popular kernels, including Gaussian kernel. Assumption 3.2 is a mild condition on the conditional density function of the error term, which is standard in quantile regression (Chen et al., 2019, 2020; Tan et al., 2022). Assumption 3.3 requires that the distribution of $\boldsymbol{X}$ have heavier tails than Gaussian to obtain standard convergence rates for the quantile regression estimates. Assumption 3.4 is commonly assumed in the sparse high-dimensional estimation literature for the support recovery (Belloni & Chernozhukov, 2011; Chen et al., 2020; Bao & Xiong, 2021). Assumption 3.5 is also a common condition in the distributed estimation literature, see also in Wang et al. (2017); Jordan et al. (2019); Chen et al. (2020). Note that $p=\mathcal{O}(N^{\nu})$, we use $\log N$ instead of the commonly used $\log(\max(p,N))$ in the convergence rates. Assumption 3.6 assumes the convergence rate and support recovery accuracy of the initial estimator, which can be satisfied by the estimator using the local sample from a single machine under Assumption 3.2-3.5 and some regularity conditions (Fan et al., 2014). We first show the convergence rate of the one-step DHSQR estimator $\widehat{\boldsymbol{\beta}}_{1,h}$.
@@
Note that $p=\mathcal{O}(N^{\nu})$ by Assumption 3.5, we have $\log(\max(N,p))\asymp\log N$, so we use $\log N$ in our theorem instead of $\log(\max(N,p))$. With a proper choice of the global and local bandwidth $h$ and $b$, we can refine the initial estimator by one iteration of our algorithm. Specifically, the convergence rate reduces from $\mathcal{O}_{\mathbb{P}}(a_{N})$ to $\mathcal{O}_{\mathbb{P}}(\max\{\sqrt{s\log N/N},\sqrt{s}a_{N}\eta\})$ with $\sqrt{s}\eta=\mathcal{O}(1)$ by Assumption 3.5. Now, we can recursively apply Theorem 3.7 to get the convergence rate of the iterative DHSQR estimator.
@@
Suppose that the initial estimator satisfies that $|\widehat{\boldsymbol{\beta}}_{0,h}-\boldsymbol{\beta}^{*}|_{2}=\mathcal{O}_{\mathbb{P}}(\sqrt{s\log N/n})$ and let $h\asymp\left(s\log N/N\right)^{1/3}$, $b\asymp\left(s\log n/n\right)^{1/3}$. For $1\leq g\leq t$, denote $\kappa(a,b;g)=\max\left\{s^{a}\left(\frac{\log n}{n}\right)^{\frac{2g+3}{6}},s^{b}\left(\frac{\log N}{n}\right)^{\frac{g+1}{2}}\right\}$ and take
$\lambda_{N,g}=C\left(\sqrt{\frac{\log N}{N}}+\kappa(\frac{5g}{6},g;g)\right),$
@@
$\left|\widehat{\boldsymbol{\beta}}_{t,h}-\boldsymbol{\beta}^{*}\right|_{2}=\mathcal{O}_{\mathbb{P}}\left(\sqrt{\frac{s\log N}{N}}+\kappa(\frac{5t+3}{6},\frac{2t+1}{2};t)\right).$ (19)
@@
Next, we provide the support recovery of the DHSQR estimator in the following two theorems. Let $\widehat{\boldsymbol{\beta}}_{t,h}=(\widehat{\beta}_{t,h}^{0},\widehat{\beta}_{t,h}^{1},\ldots,\widehat{\beta}_{t,h}^{p})$ and $\widehat{S}_{t}=\{j:\widehat{\beta}_{t,h}^{p}\neq 0,j\in\mathbb{N}_{+}\}$ be the support of $\widehat{\boldsymbol{\beta}}_{t,h}$, where $t\geq 1$.
@@
Based on Theorem 3.9, we can show a weaker beta-min condition for the support recovery result of the iterative DHSQR estimator.

### v3

@@ 2.3 Distributed Estimation with a Double-smoothing Shifted Loss Function
Suppose the random samples ${\cal Z}^{N}=\{(\boldsymbol{X}_{i},Y_{i})\}_{i=1}^{N}$ are randomly stored in $m$ machines $\mathcal{M}_{1},\ldots,\mathcal{M}_{m}$ with the equal local sample size that $n=N/m$. Without loss of generality, we assume that $\mathcal{M}_{1}$ is the central machine and denote those samples in the $k$-th machine as $\{(\boldsymbol{X}_{i},Y_{i})\}_{i\in\mathcal{M}_{k}}$ with $|\mathcal{M}_{k}|=n$, for $k=1,\ldots,m$. Based on the initial estimator $\widehat{\boldsymbol{\beta}}_{0}$, every local machines can compute the transformed samples as $\{(\widetilde{\boldsymbol{X}}_{i,h},\widetilde{Y}_{i,h})\}_{i\in\mathcal{M}_{k}}$ according to (6) and (7). For ease of notation, let
@@
$\widehat{\boldsymbol{D}}_{h}=\frac{1}{m}\sum_{k=1}^{m}\widehat{\boldsymbol{D}}_{k,h}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}},$
@@
According to the Taylor expansion of $\mathcal{L}_{N}(\boldsymbol{\beta})$ around $\widehat{\boldsymbol{\beta}}_{0}$, we have
@@
It is easy to see that $\partial\mathcal{L}_{N}(\widehat{\boldsymbol{\beta}}_{0})$ and $\widehat{\boldsymbol{D}}_{h}$ can be simply calculated by averaging the local ones. However, the burden of transmitting the local $(p+1)\times(p+1)$ covariance matrix $\widehat{\boldsymbol{D}}_{k,h}$ is heavy when $p$ is large. To save the communication cost, we replace the global Hessian $\widehat{\boldsymbol{D}}_{h}$ with the local Hessian $\widehat{\boldsymbol{D}}_{1,b}$. Here, $h$ and $b$ denote the global bandwidth and local bandwidth, respectively, and we assume $b\geq h\geq 0$. Thus we can rewrite (11) as
$\mathcal{L}_{N}(\boldsymbol{\beta},\widehat{\boldsymbol{D}}_{h})$ $=\underbrace{\mathcal{L}_{N}(\boldsymbol{\beta},\widehat{\boldsymbol{D}}_{1,b})}_{(i)\text{ Shifted loss}}$ (12)
$+\underbrace{\mathcal{O}_{\mathbb{P}}\left\{\|\widehat{\boldsymbol{D}}_{h}-\widehat{\boldsymbol{D}}_{1,b}\|_{op}\cdot|\boldsymbol{\beta}-\widehat{\boldsymbol{\beta}}_{0}|_{2}^{2}\right\}}_{(ii)\text{ Approximation error}}.$
The second term in (12) is from the Cauchy–Schwarz inequality. Note that the substituted local Hessian matrix is flexibly controlled by a local bandwidth $b$ instead of the global bandwidth $h$, which ensures that $\|\widehat{\boldsymbol{D}}_{h}-\widehat{\boldsymbol{D}}_{1,b}\|_{op}=o_{\mathbb{P}}(1)$ (Detailed proof can be referred in Appendix A.3). Remove the terms that are independent of $\boldsymbol{\beta}$ in (i) and the negligible approximation error (ii) in (12), the shifted loss function can be simplified to
@@
where $z_{N}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{Y}_{i,h}$. Up to now, we only need to focus on the shifted loss function in (13) instead of the pseudo global loss function in (10) for higher communication efficiency. Specifically, we define the one-step distributed estimator as
@@
Note that the local machines only need to compute two vectors $z_{n,k}=\frac{1}{n}\sum_{i\in\mathcal{M}_{k}}\widetilde{\boldsymbol{X}}_{i,h}\widetilde{Y}_{i,h}$ and $\widehat{\boldsymbol{D}}_{k,h}\widehat{\boldsymbol{\beta}}_{0}=\frac{1}{n}\sum_{i\in\mathcal{M}_{k}}\widetilde{\boldsymbol{X}}_{i,h}(\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{0})$ and then broadcast them to the central machine with communication cost of $\mathcal{O}(mp)$. There is no need to communicate the $(p+1)\times(p+1)$ covariance matrix $\widehat{\boldsymbol{D}}_{k,h}$. The central machine first calculates $z_{N}$ and $\widehat{\boldsymbol{D}}_{h}\widehat{\boldsymbol{\beta}}_{0}$ by simple averaging, then solves (14) via some well-learned algorithms, e.g., the PSSsp algorithm (Schmidt, 2010), the active set algorithm (Solntsev et al., 2015) and the coordinate descent algorithm (Wright, 2015).
Given $\widehat{\boldsymbol{\beta}}_{1,h}$ as the estimator from the first iteration, we can similarly construct an iterative distributed estimation procedure. Specifically, in the $t$-th iteration, we update the pseudo covariates and responses in (6) and (7) by substituting $\widehat{\boldsymbol{\beta}}_{0}$ with $\widehat{\boldsymbol{\beta}}_{t-1,h}$,
@@
for $i=1,\ldots,N$, where $\widehat{e}^{(t)}_{i,h}=Y_{i}-\boldsymbol{X}_{i}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{t,h}$. Similar with (14), the distributed estimator in the $t$-th iteration is
@@
Algorithm 1 Distributed high-dimensional sparse quantile regression (DHSQR).
@@
$(\widehat{\boldsymbol{D}}_{1,b}^{(t)}-\widehat{\boldsymbol{D}}_{h}^{(t)})\widehat{\boldsymbol{\beta}}_{t-1,h}=\widehat{\boldsymbol{D}}_{1,b}^{(t)}\widehat{\boldsymbol{\beta}}_{t-1,h}-\frac{1}{m}\sum_{k=1}^{m}\widehat{\boldsymbol{D}}_{k,h}^{(t)}\widehat{\boldsymbol{\beta}}_{t-1,h},\quad\boldsymbol{z}_{N}^{(t)}=\frac{1}{m}\sum_{k=1}^{m}\boldsymbol{z}_{n,k}^{(t)}.$
@@
In this paper, we adopt the coordinate descent algorithm to solve (17). For the choice of the initial estimator $\widehat{\boldsymbol{\beta}}_{0}$, we take the solution of $\ell_{1}$-penalized QR regression using the local data on the central machine, which can be solved by the R package ”quantreg(Koenker, 2005) or ”conquer(He et al., 2023). Other types of initialization are also available as long as they satisfy Assumption 3.6 in Section 3. We summarize the entire distributed estimation procedure in Algorithm 1.
Space complexity. In each local machine $\mathcal{M}_{i}$, DHSQR method necessitates storing $\{\boldsymbol{X}_{i}\}_{i\in\mathcal{M}_{i}}\in\mathbb{R}^{(p+1)\times n}$, $\{Y_{i}\}_{i\in\mathcal{M}_{i}}\in\mathbb{R}^{n}$, $\{\tilde{\boldsymbol{X}}_{i,h}^{(t)}\}_{i\in\mathcal{M}_{i}}\in\mathbb{R}^{(p+1)\times n}$, and $\{\tilde{Y}_{i,h}^{(t)}\}_{i\in\mathcal{M}_{i}}\in\mathbb{R}^{n}$, resulting in a space complexity of order $\mathcal{O}(np)$. Additionally, to solve the lasso problem, we require storing a $(p+1)\times(p+1)$ matrix, resulting in a space complexity not exceeding $\mathcal{O}\left(p^{2}\right)$. Hence, the overall space complexity is of order $\mathcal{O}\left(np+p^{2}\right)$ for each local machine. The total space complexity of the total system sums up to $\mathcal{O}\left(Np+mp^{2}\right)$.
Remark 2.2.
Note that the assumption that the samples are randomly and even stored across the local machines is commonly required in literature (Jordan et al., 2019; Fan et al., 2021). It is worthy pointing out that the proposed algorithm is still effective if the samples in other local machines are not randomly distributed as long as the subsample on the first machine (central machine) is randomly selected from the entire sample, which is also claimed in Remark 1 in Chen et al. (2020).
@@ 3 Statistical Guarantees
$S=\{j:\beta_{j}\neq 0,j\in\mathbb{N}_{+}\}\subseteq\{1,\ldots,p\},$
@@
Assume that the kernel function $K(\cdot)$ is symmetric, non-negative, bound and integrates to one. In addition, the kernel function satisfies that $\int_{-\infty}^{\infty}u^{2}K(u)\mathrm{d}u<\infty$ and $\min_{|u|\leq 1}K(u)>0$. We further assume $K(\cdot)$ is differentiable and its derivative $K^{\prime}(\cdot)$ is bounded.
@@
There exist $f_{2}\geq f_{1}>0$ such that $f_{1}\leqslant f_{\varepsilon\mid\boldsymbol{X}}(0)\leqslant f_{2}$ almost surely (for all $\boldsymbol{X}$). Moreover, there exists some $l_{0}$ such that
$\left|f_{\varepsilon\mid\boldsymbol{X}}(u)-f_{\varepsilon\mid\boldsymbol{X}}(v)\right|\leq l_{0}|u-v|,$
for any $u,v\in\mathbb{R}$ and all $\boldsymbol{X}$.
@@
The random covariate $\boldsymbol{X}\in\mathbb{R}^{p}$ is sub-Gaussian: there exists some $c_{1}>0$ such that
${\mathbb{P}}\left(\left|\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\Sigma}^{-1/2}\delta\right|\geqslant c_{1}t\right)\leq 2e^{-t^{2}/2},$
for every unit vector $\delta$ and $t>0$, where $\boldsymbol{\Sigma}=\mathbb{E}(\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}})$.
@@
Denote $\boldsymbol{I}=\mathbb{E}\left\{f_{\varepsilon\mid\boldsymbol{X}}(0)\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}\right\}$, then $\boldsymbol{I}$ satisfies that
$\left\|\boldsymbol{I}_{S^{c}\times S}\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\leq 1-\alpha,$
for some $10<\alpha<$ 1. Moreover, we assume that $\lambda_{-}\leq\Lambda_{\min}(\boldsymbol{I})\leq\Lambda_{\max}(\boldsymbol{I})\leq\lambda_{+}$ for some $\lambda_{-},\lambda_{+}>0$.
@@
The dimension $p$ satisfies $p=\mathcal{O}\left(N^{\nu}\right)$ for some $\nu>0$. The local sample size $n$ satisfies $n\geq N^{c}$ for some $10<c<$ 1, and the sparsity level $s$ satisfies $s=\mathcal{O}\left(n^{r}\right)$ for some $30\leq r<1/$ 3.
@@
Assumption 3.1 imposes some regularity conditions on the kernel function $K(\cdot)$, which is satisfied by many popular kernels, including the Gaussian kernel. Assumption 3.2 is a mild condition on the smoothness of the conditional density function of the error term, which is standard in quantile regression (Chen et al., 2019, 2020; Tan et al., 2022). Assumption 3.3 requires that the distribution of $\boldsymbol{X}$ have heavier tails than Gaussian to obtain standard convergence rates for the quantile regression estimates. Assumption 3.4 is known as the irrepresentable condition, which is commonly assumed in the sparse high-dimensional estimation literature for the support recovery (Belloni & Chernozhukov, 2011; Chen et al., 2020; Bao & Xiong, 2021). Assumption 3.5 is also a common condition in the distributed estimation literature, see also in Wang et al. (2017); Jordan et al. (2019); Chen et al. (2020). Note that $p=\mathcal{O}(N^{\nu})$, we use $\log N$ instead of the commonly used $\log(\max(p,N))$ in the convergence rates. Assumption 3.6 assumes the convergence rate and support recovery accuracy of the initial estimator, which can be satisfied by the estimator using the local sample from a single machine under Assumption 3.2-3.5 and some regularity conditions (Fan et al., 2014). We first show the convergence rate of the one-step DHSQR estimator $\widehat{\boldsymbol{\beta}}_{1,h}$.
@@
With a proper choice of the global and local bandwidth $h$ and $b$, we can refine the initial estimator by one iteration of our algorithm. Specifically, the convergence rate reduces from $\mathcal{O}_{\mathbb{P}}(a_{N})$ to $\mathcal{O}_{\mathbb{P}}(\max\{\sqrt{s\log N/N},\sqrt{s}a_{N}\eta\})$ with $\sqrt{s}\eta=o(1)$ by Assumption 3.5. Now, we can recursively apply Theorem 3.7 to get the convergence rate of the iterative DHSQR estimator. For $1\leq g\leq t$, we denote
$\kappa(a,b,g)=\max\left\{s^{a}\left(\frac{\log n}{n}\right)^{\frac{2g+3}{6}},s^{b}\left(\frac{\log N}{n}\right)^{\frac{g+1}{2}}\right\}.$
@@
Suppose that the initial estimator satisfies that $|\widehat{\boldsymbol{\beta}}_{0,h}-\boldsymbol{\beta}^{*}|_{2}=\mathcal{O}_{\mathbb{P}}(\sqrt{s\log N/n})$ and let $h\asymp\left(s\log N/N\right)^{1/3}$, $b\asymp\left(s\log n/n\right)^{1/3}$. For $1\leq g\leq t$, take
$\lambda_{N,g}=C\left(\sqrt{\frac{\log N}{N}}+\kappa(\frac{5g}{6},g,g)\right),$
@@
$\left|\widehat{\boldsymbol{\beta}}_{t,h}-\boldsymbol{\beta}^{*}\right|_{2}=\mathcal{O}_{\mathbb{P}}\left(\sqrt{\frac{s\log N}{N}}+\kappa(\frac{5t+3}{6},\frac{2t+1}{2},t)\right).$ (19)
@@
Next, we provide the support recovery of the one-step and $t$-th iteration DHSQR estimators in the following two theorems. Let $\widehat{\boldsymbol{\beta}}_{t,h}=(\widehat{\beta}_{t,h}^{0},\widehat{\beta}_{t,h}^{1},\ldots,\widehat{\beta}_{t,h}^{p})$ and
$\widehat{S}_{t}=\{j:\widehat{\beta}_{t,h}^{p}\neq 0,j\in\mathbb{N}_{+}\},$
be the support of $\widehat{\boldsymbol{\beta}}_{t,h}$, where $t\geq 1$.
@@
Based on Theorem 3.9, we can show a weaker beta-min condition for the support recovery result of the $t$-th iteration DHSQR estimator.